OpenStudy (ksaimouli):

solve for x cos^2x-sin^2x=sin x

OpenStudy (ksaimouli):

\[\cos^2x-\sin^2x =sinx\]

OpenStudy (eyust707):

I see trig... I see trig stuff squared... im thinkin trig identities right off the bat!

terenzreignz (terenzreignz):

cox 2x comes to mind...

terenzreignz (terenzreignz):

but it's useless, I think?

OpenStudy (eyust707):

no i like the thinking!

terenzreignz (terenzreignz):

LOL, this is a quadratic equation in disguise

terenzreignz (terenzreignz):

cos^2 x = 1 - sin^2 x right? so we can turn it into... 1 - sin^2 x - sin^2 x = sin x

OpenStudy (anonymous):

\[\cos^2x-\sin^2=1-2\sin^2x\]

OpenStudy (anonymous):

oops :P ditto

terenzreignz (terenzreignz):

Oh, you can do it from here? great :)

OpenStudy (anonymous):

I'm not the asker @ksaimouli can you take it from here?

terenzreignz (terenzreignz):

derp, I forgot about that (sorry)